OpenStudy (anonymous):

Consider the following reaction: 2CH3OH ---> 2CH4+O2 delta H= +252.8kJ Calculate the amount of heat transferred when 22.0g of CH3OH(g) is decomposed by this reaction at constant pressure. Include units

OpenStudy (matt101):

This is an endothermic reaction because the enthalpy (ΔH) is positive. You can also see from the reaction you provided that for every 2 moles of CH3OH used up, 252.8 kJ of energy is used up too. That means for 1 mole of CH3OH, half as much energy would be used up, or 126.4 kJ. We just need to find how many moles (n) of CH3OH we have, and we do that using its molar mass. n = (22 g)/(32 g/mol) = 0.6875 mol Now set up a ratio - if it takes 126.4 kJ to decompose 1 mole of CH3OH, how much energy does it take to decompose 0.6875 moles? x/(0.6875 mol) = (126.4 kJ)/(1 mol) --> x = 86.9 kJ